solve : cot(3 + sin(x)) =1
Answer by Alan3354(69443)

solve : cot(3 + sin(x)) =1
3 + sin(x) = pi/4 + n*pi radians, n = integer
sin(x) = pi/4 - 3 = -2.21 - No real number solution
-----
sin(x) = pi/4 + pi - 3
sin(x) = ~ 0.926
x = 1.186 radians = ~ 67.97 degs
=================
sin(x) = pi/4 + 2pi - 3
sin(x) = ~4.06 - No real # solution